# Non-equilibrium ionization in the multiphase circumgalactic medium—impact on quasar absorption-line analyses

### v1

@@ Abstract
This paper presents an updated framework for studying the ionizing conditions and elemental abundances of photoionized, metal-enriched quasar absorption systems. The standard assumption of ionization equilibrium invoked in absorption line analyses requires gas to cool on longer timescales than ionic recombination (). However, this assumption may not be valid at high metallicities due to enhanced cooling losses. This work presents a suite of time-dependent photoionization (TDP) models that self-consistently solve for the ionization state of rapidly cooling gas irradiated by the extragalactic ultraviolet background (UVB). The updated framework explores various revised UVBs from recent studies, a range of initial temperatures, and different elemental abundance patterns to quantify the effects of TDP on the observed ion fractions. A metal-enriched () C IV absorption system at previously studied using photoionization equilibrium (PIE) models is re-examined under the TDP framework. The main findings are as follows: (1) varying prescriptions for the underlying UVB or adopting initial temperatures change TDP ion fractions by up to a factor of three, but the adopted relative elemental abundance pattern affects ion fractions by at most 40%; (2) the inferred gas densities are consistent between PIE and TDP, but under TDP solar metallicity cannot be ruled out at more than 2- significance and a modest non-solar [C/] is needed to explain the observed relative ion abundances. Extending the TDP analyses to a larger sample of super-solar absorption components with high signal-to-noise absorption spectra is needed to quantify the fraction of metal absorbers originating in rapid cooling gas.
@@
Both photo- and collisional ionization models typically assume equilibrium conditions. Under this assumption, the gas must cool slowly relative to ionic recombination timescales, so a steady state can be assumed. However, at high metallicities, cooling losses are enhanced through collisional, free-bound, and free-free interactions; the cooling time becomes shorter than ionic recombination, resulting in a departure from ionization equilibrium (see e.g., Gnat & Sternberg, 2007). In such cases, a steady state cannot be assumed, and equations of ionization balance must be solved simultaneously with energy conservation to compute time-dependent ion fractions.
@@
Studies of hot gas have considered both collisional equilibrium and non-equilibrium models (see e.g., Tripp et al., 2011; Meiring et al., 2013; Hussain et al., 2015; Pachat et al., 2017; Nevalainen et al., 2017; Rosenwasser et al., 2018). The cool phase has been established to be in photoionization equilibrium (see e.g., Zahedy et al., 2019; Qu et al., 2022; Sameer et al., 2021; Zahedy et al., 2021; Cooper et al., 2021), justified by the inferred sub-solar metallicities which are not sufficient to cause rapid cooling. Photoionization equilibrium models have also been used for the warm-hot phase (see e.g., Sankar et al., 2020; Kumar et al., 2024), but the inferred metallicities at a solar level or above pose a challenge because of enhanced cooling losses expected at these enrichment levels. Additionally, Kumar et al. (2024) find a discrepancy between the ionization equilibrium model temperature and line width temperature in the C IV-selected warm-hot CGM. These caveats motivate the usage of non-equilibrium photoionization models for metal-enriched warm-hot gas.
@@
Non-equilibrium (or time-dependent) photoionization by the extragalactic ultraviolet background (UVB) has been discussed by Oppenheimer & Schaye (2013) and Gnat (2017). Both studies demonstrate that the ionization state of a rapidly cooling gas differs from the equilibrium scenario. In addition, photoionization modifies the initial overabundance of high-ionization species arising from recombination lags, further amplifying the discrepant ionization balance. An important distinction between photoionization equilibrium and non-equilibrium is the added dependence of ion ratios on gas metallicity and temperature. Therefore, systematic differences in inferred gas properties may exist using photoionization equilibrium versus non-equilibrium models. However, before comparing non-equilibrium models with absorption line measurements, some systematics regarding this class of models are worth highlighting.
@@
The primary ingredient for photoionization models is the assumed ionizing radiation background. The CGM, being spatially extended, remains largely unaffected by ionizing radiation arising locally within the galaxy halo (e.g., Qu et al., 2023). Instead, the CGM is ionized by the UVB, which has contributions from quasars and star-forming galaxies. There are various prescriptions available for the ionizing background (see e.g., Haardt & Madau, 2001, 2012; Khaire & Srianand, 2019; Faucher-Giguère, 2020); these prescriptions differ in their amplitude and slope, influencing ion fractions in photoionized gas (see e.g., Chen et al., 2017; Zahedy et al., 2019; Lehner et al., 2022). Non-equilibrium models from Oppenheimer & Schaye (2013) and Gnat (2017) use UVB prescriptions from Haardt & Madau (2001) and Haardt & Madau (2012, HM12). In the past decade, extensive efforts have been directed towards revising UVB models to incorporate updated statistics of Ly forest absorbers and improved constraints on AGN luminosity functions/x-ray extragalactic background. As such, recent prescriptions from (Khaire & Srianand, 2019, KS19) and (Faucher-Giguère, 2020, FG20) have not been incorporated into time-dependent photoionization models. In addition to the ionizing background, the adopted initial temperature from which the gas is assumed to cool is another important ingredient in non-equilibrium ionization models. This is because the initial temperature sets the starting ionic composition of the gas, which affects the subsequent ionization state of the gas through time-dependent cooling. Finally, the adopted elemental abundance pattern sets the relative cooling losses from individual elements (see e.g., Gnat & Ferland, 2012), which are coupled to the ion fractions. It is crucial to explore the variance in non-equilibrium ion fractions resulting from uncertainties in these three model assumptions.
@@
The additional dependence of ion fractions on gas temperature and metallicity under non-equilibrium photoionization is important to address when comparing model predictions with absorption line measurements. In absorption line systems with ions of sufficiently different mass, the line broadening can be decomposed into thermal and non-thermal contributions (see e.g., Rauch et al., 1996; Rudie et al., 2019; Zahedy et al., 2019; Qu et al., 2022; Kumar et al., 2024). In situations where the gas temperature can be determined using the observed absorption line widths, the assumed temperature in the photoionization calculations can be fixed to the line width temperature to reduce degeneracies between different gas properties during ionization modeling (see e.g., Sameer et al., 2024).
@@
This paper is organized as follows. In § 2, the construction of ionization models is presented. In § 3, the impact of varying individual model assumptions on non-equilibrium fractions is discussed, followed by comparing model predictions with absorption line measurements. In § 4, a summary of the key findings from the study is presented alongside caveats.
2 Method
This section presents the details of building ionization models for examining absorption line measurements. Specifically in § 2.1, key ingredients for photoionization equilibrium models and caveats of using such models are highlighted, and in § 2.2, time-dependent photoionization as an alternative to equilibrium models for examining metal-enriched gas is presented.
2.1 Photoionization equilibrium
Ionization analyses of cool and warm-hot diffuse gases typically assume photo-ionization equilibrium (PIE). The gas is irradiated by the extragalactic UVB. Some available UVB prescriptions are shown in Figure 1. The ionization state of gas in PIE is set by a balance between photoionization and ionic recombination. A useful quantity for characterizing PIE fractions is the ionization parameter , which quantifies the ratio of ionizing photons to baryons. PIE fractions primarily depend on the ionization parameter and are only weakly dependent on metallicity and temperature. Therefore, measured ion ratios from absorption line systems can be compared with expectations from PIE to constrain the ionization parameter robustly. Under the assumption of a fixed radiation field, a constraint on the ionization parameter translates to a measure of the gas density.
@@ 2.1 Photoionization equilibrium
A test for the validity of PIE is comparing the ionization model temperature with the line width temperature. PIE assumes thermal equilibrium, with the gas temperature set by a balance between radiative cooling from collisional processes and photoheating by the UVB. Figure 2 depicts the cooling and heating rates for gas at solar and sub-solar metallicities, assuming a fiducial gas density of . The equilibrium temperature depends on the gas metallicity, with higher enrichment levels leading to lower temperatures. The equilibrium temperature corresponding to the best-fit ionization model can be compared with the temperature estimated by comparing the absorption line widths from ionic species with sufficiently different masses (see e.g., Qu et al., 2022; Kumar et al., 2024).
A recent investigation of C IV absorbers at from Kumar et al. (2024, hereafter Paper I) pursued ionization modeling of 12 kinematically resolved absorption components. The equilibrium temperature for each component was determined based on the best-fit gas density and metallicity. The kinematic temperature was obtained using the best fit line widths of the H I and C IV components that were found to occupy the same phase during ionization modeling. The line width temperature was found to exceed the PIE temperature at a >3- level for the C IV-bearing gas phase at (see §3.2). Incidentally, this phase also had the highest inferred metallicity in the sample, with . The cooling time for this system is 7.5 Myr, shorter than the C V C IV recombination time of 12 Myr. The mismatched cooling and recombination time scales indicate that the gas is likely out of equilibrium in this metal-enriched system.
In principle, the temperature discrepancy can be addressed by fixing the gas temperature adopted for PIE calculations to be consistent with the observed thermal broadening (see e.g., Hussain et al., 2015; Pachat et al., 2017). Physically, such a model assumes the presence of an unknown heating source, elevating the equilibrium temperature compared to UVB-only heating (Figure 3). In the measured temperature range , PIE ion fractions show only modest variations (see e.g., Gnat, 2017, also Figure 4). Therefore, the gas properties inferred from this modified calculation would be consistent with the original PIE results presented in Paper I (summarized in § 3.2). Although this approach may resolve the tension between the PIE temperature and the observed line broadening, caveats remain in identifying a suitable astrophysical source that can provide heating without modifying the spectral shape of the ionizing radiation (see § 4). As discussed in the next section and in § 3.2, time-dependent photoionization (TDP) offers an alternative framework for reproducing absorption line measurements by modeling rapid gas cooling at high metallicities.
2.2 Time-dependent photoionization
@@
Chemically enriched gas clouds are expected to experience enhanced cooling losses through collisional, free-bound, and free-free interactions. In ionization equilibrium, because the gas cools slowly, ion fractions can be assumed to be in a steady state at all times (). However, a steady state cannot be assumed if cooling is rapid. In such cases, equations of ionization balance must be solved simultaneously with energy conservation to specify the ionization state of the gas over time. At high metallicities (), when cooling is enhanced, the gas becomes "overionized" relative to the equilibrium scenario because of an overabundance of high ions caused by recombination lag (see e.g., Gnat & Sternberg, 2007).
@@
Non-equilibrium collisional ionization models from Gnat & Sternberg (2007) have been compared with absorption line systems arising from hot gas (see e.g., Tripp et al., 2011; Meiring et al., 2013; Hussain et al., 2015; Nevalainen et al., 2017; Rosenwasser et al., 2018). However, photoionization from the extragalactic UVB cannot be neglected at temperatures relevant for some high-ionization species. The inclusion of a radiation background in time-dependent calculations modifies the ionization state of gas through photoionization and photoheating (see e.g., Oppenheimer & Schaye, 2013; Gnat, 2017). This is illustrated in the left panel of Figure 3, which shows the differences in the predicted ion fractions between PIE and TDP models for three commonly adopted tracers of warm-hot ionized gas, C IV, N V, and O VI, in QSO absorption-line studies.
Note that the PIE fractions vary with gas density but do not depend on metallicity, while the predicted TDP fractions have a simultaneous density and metallicity dependence. The discrepancy between TDP and PIE is therefore metallicity dependent. For a solar metallicity gas of density , O VI can be overabundant by an order of magnitude at . N V behaves similarly as O VI, but C IV is underabundant relative to PIE at this density because of an overabundance in C V at these cooler temperatures.
@@
The photoionization models presented in this paper are calculated using an updated CLOUDY photoionization code (version 22.01, see Ferland et al., 2017; Chatzikos et al., 2023). In particular, the hazy1 documentation provides a non-equilibrium cooling calculation script for modeling the radiative time-dependent cooling in a plane parallel gas slab without any ionizing radiation, following Gnat & Sternberg (2007). The default script is modified to include ionizing radiation from a UVB 111See this ReadMe for a tutorial on building a suite of TDP models.. The cooling is assumed to be isochoric, and the calculations stop at thermal equilibrium or , whichever condition is achieved first. The gas is also assumed to be optically thin for these calculations.
@@
It was shown in Paper I that under the PIE assumption C IV absorption preferentially arise in diffuse gas of and high metallicities of . TDP fractions in this density-metallicity range can depart significantly from the equilibrium values. Because gas properties are inferred by comparing measured ion ratios with ionization model predictions, the results of such a comparison can differ between TDP and PIE.
This work will re-examine the highest metallicity C IV component from Paper I under a TDP framework. To do so, the TDP calculations highlighted before are performed for a grid of gas densities and metallicities. An ionization analysis comparing absorption line measurements with TDP models reveals how inferred gas properties change compared to a PIE analysis. Note that for the systems of interest, the gas temperature is well-constrained from available absorption line widths of elements with sufficiently different masses. In fact, the temperature estimate using absorption line widths was found to be discrepant with the PIE model temperature. Therefore, when comparing TDP models with absorption line measurements, the model temperature is fixed to the line width temperature. A direct implication of this approach is that the gas is in a transient phase, still undergoing cooling and not yet having reached thermal equilibrium. This is distinct from the steady-state PIE solution discussed in the previous section which requires additional heating from some unspecified source (see § 4 for further discussion).
3 Analysis
Time-dependent photoionization models provide a self-consistent framework for examining metal-enriched absorption systems. However, the underlying assumptions for TDP models can have a notable effect on the expected ion fractions. Here the effect of varying individual ingredients in the TDP models is evaluated in § 3.1 and a comparison between the TDP model predictions and empirical measurements of a high-metallicity C IV absorber is presented in § 3.2.
@@ 3.1 Model assumptions
Three key ingredients of time-dependent photoionization models are the underlying ionizing radiation, the initial temperature from which cooling occurs and the adopted abundance pattern. Reasonable choices for each ingredient are explored in the subsections here with § 3.1.1 discussing the impact of varying the adopted UVB, § 3.1.2 evaluating the effect of the chosen initial temperature, and § 3.1.3 examining the impact of using non-solar abundance patterns.
@@ 3.1.1 UVB prescription
The right panel in Figure 4 shows the predicted TDP fractions for C IV, N V, and O VI for three different UVBs at , including HM12, KS19, and FG20. The predictions are obtained using the procedure described in § 2.2. The high-temperature () peak comes from collisional ionization, while the low-temperature peak occurs because of photoionization. The fiducial choice of and for the density and metallicity are relevant to enriched C IV absorbers (§ 3.2). While ion fractions for C IV are robust against UVB choices, N V and O VI TDP fractions can differ by a factor of up to three at . To facilitate a more focused discussion on the effect of TDP, FG20 is adopted as the fiducial UVB spectrum in the subsequent discussions.

### v2

@@ Abstract
This paper presents an updated framework for studying the ionizing conditions and elemental abundances of photoionized, metal-enriched quasar absorption systems. The standard assumption of ionization equilibrium invoked in absorption line analyses requires gas to cool on longer timescales than ionic recombination (). However, this assumption may not be valid at high metallicities due to enhanced cooling losses. This work presents a suite of time-dependent photoionization (TDP) models that self-consistently solve for the ionization state of rapidly cooling gas irradiated by the extragalactic ultraviolet background (UVB). The updated framework explores various revised UVBs from recent studies, a range of initial temperatures, and different elemental abundance patterns to quantify the effects of TDP on the observed ion fractions. A metal-enriched () C IV absorption system at previously studied using photoionization equilibrium (PIE) models is re-examined under the TDP framework. The main findings are as follows: (1) varying prescriptions for the underlying UVB or adopting initial temperatures (with the starting ionization state in collisional equilibrium) change TDP ion fractions by up to a factor of three and ten respectively, but the adopted relative elemental abundance pattern affects ion fractions by at most 40%; (2) the inferred gas densities are consistent between PIE and TDP, but under TDP solar metallicity cannot be ruled out at more than 2- significance and a non-solar [C/] is robustly constrained from the observed relative ion abundances. Extending the TDP analyses to a larger sample of super-solar absorption components with high signal-to-noise absorption spectra is needed to quantify the fraction of metal absorbers originating in rapid cooling gas.
@@
Both photo- and collisional ionization models typically assume equilibrium conditions. Under this assumption, the gas must cool slowly relative to ionic recombination timescales, so a steady state can be obtained. However, at high metallicities, cooling losses are enhanced through collisional, free-bound, and free-free interactions; the cooling time becomes shorter than ionic recombination, resulting in a departure from ionization equilibrium (see e.g., Gnat & Sternberg, 2007). In such cases, a steady state cannot be assumed, and equations of ionization balance must be solved simultaneously with energy loss to compute time-dependent ion fractions.
@@
Studies of hot gas have considered both collisional equilibrium and non-equilibrium models (see e.g., Tripp et al., 2011; Meiring et al., 2013; Hussain et al., 2015; Pachat et al., 2017; Nevalainen et al., 2017; Rosenwasser et al., 2018). The cool phase has been established to be in photoionization equilibrium (see e.g., Zahedy et al., 2019; Qu et al., 2022; Sameer et al., 2021; Zahedy et al., 2021; Cooper et al., 2021), justified by the inferred sub-solar metallicities which lead to slow net cooling. Photoionization equilibrium (PIE) models have also been used for the warm-hot phase (see e.g., Sankar et al., 2020; Kumar et al., 2024), but the inferred metallicities at a solar level or above pose a challenge because of enhanced cooling losses expected at these enrichment levels. Additionally, Kumar et al. (2024) find a discrepancy between the ionization model temperature and the observed thermal broadening in C IV absorbers after accounting for non-thermal motions in the gas. These caveats motivate the usage of non-equilibrium photoionization models for metal-enriched warm-hot gas.
@@
Non-equilibrium (or time-dependent) photoionization by the extragalactic ultraviolet background (UVB) has been discussed by Oppenheimer & Schaye (2013) and Gnat (2017). The observed ionization state of a gas out of equilibrium can depart significantly from a gas in equilibrium, assuming otherwise identical physical properties (density, temperature, and elemental abundances). Therefore, systematic differences in inferred gas properties may exist using photoionization equilibrium versus non-equilibrium models. However, before comparing non-equilibrium models with absorption line measurements, some systematics regarding this class of models are worth highlighting.
@@
The primary ingredient for photoionization models is the assumed ionizing radiation background. The CGM, being spatially extended, remains largely unaffected by ionizing radiation arising locally within the galaxy halo (e.g., Qu et al., 2023). Instead, the CGM is ionized by the UVB, which has contributions from quasars and star-forming galaxies. There are various prescriptions available for the ionizing background (see e.g., Haardt & Madau, 2001, 2012; Khaire & Srianand, 2019; Faucher-Giguère, 2020); these prescriptions differ in their amplitude and slope, influencing ion fractions in photoionized gas (see e.g., Chen et al., 2017; Zahedy et al., 2019; Lehner et al., 2022). Non-equilibrium models from Oppenheimer & Schaye (2013) and Gnat (2017) use UVB prescriptions from Haardt & Madau (2001) and Haardt & Madau (2012, HM12). In the past decade, extensive efforts have been directed towards revising UVB models to incorporate updated statistics of Ly forest absorbers and improved constraints on AGN luminosity functions/x-ray extragalactic background. As such, recent prescriptions from (Khaire & Srianand, 2019, KS19) and (Faucher-Giguère, 2020, FG20) have not been incorporated into time-dependent photoionization (TDP) models. In addition to the ionizing background, the adopted initial temperature from which the gas is assumed to cool is another important ingredient in non-equilibrium ionization models. This is because the initial temperature sets the starting ionic composition of the gas, which affects the subsequent ionization state of the gas through time-dependent cooling. Finally, the adopted elemental abundance pattern sets the relative cooling losses from individual elements (see e.g., Gnat & Ferland, 2012), which are coupled to the ion fractions. It is crucial to explore the variance in non-equilibrium ion fractions resulting from uncertainties in these three model assumptions.
@@
The additional dependence of ion fractions on gas temperature and metallicity under non-equilibrium photoionization is important to address when comparing model predictions with absorption line measurements. In absorption line systems with species of sufficiently different mass and similar ionization stage, the line broadening can be decomposed into thermal and non-thermal contributions (see e.g., Rauch et al., 1996; Rudie et al., 2019; Zahedy et al., 2019; Qu et al., 2022; Kumar et al., 2024). In situations where the gas temperature can be determined using the observed absorption line widths, the assumed temperature in the photoionization calculations can be fixed to the line width temperature to reduce degeneracies between different gas properties during ionization modeling (see e.g., Sameer et al., 2024).
@@
This paper is organized as follows. In § 2, the construction of ionization models are presented, alongside the impact of varying individual model assumptions on non-equilibrium fractions. In § 3, model predictions are compared with absorption line measurements. In § 4, a summary of the key findings from the study is presented alongside caveats.
2 Model analysis
Recall that in Kumar et al. (2024), the authors analyzed a sample of C IV-selected absorbers identified at in high-resolution QSO spectra and derived constraints for the density of the absorbing gas based on a suite of ionization transitions. The high spectral resolving power enabled the authors to decompose the observed line widths into thermal and non-thermal components, placing strong constraints on the gas temperature in the range of – and turbulent velocities of . By comparing observed relative ionic abundances of kinematically matched absorbing components with the expectations from photoionization equilibrium models, they showed that C IV absorption primarily traces diffuse gas of density – that is enriched with metallicity of (with oxygen, silicon, and neon as tracers of elements) but with an enhanced carbon abundance of . However, this analysis also uncovered two C IV components for which super solar metallicities () and a cooler gas temperature than inferred from the observed line profiles are found (e.g., component 3 of the absorber at in Kumar et al., 2024). To explore the possibility of the gas being out of equilibrium due to rapid cooling, here we examine the distinct features associated with equilibrium and non-equilibrium photoionization models and discuss their associated caveats.
2.1 Photoionization equilibrium (PIE)
The photoionization models presented in this paper are calculated using an updated CLOUDY photoionization code (version 22.01, see Ferland et al., 2017; Chatzikos et al., 2023) for a plane-parallel slab of gas irradiated by the extragalactic UVB. Some available UVB prescriptions are shown in Figure 1. The ionization state of gas in PIE is set by a balance between photoionization and ionic recombination. The equilibrium calculations are performed over a grid of temperatures while including ionizing radiation from the extragalactic UVB, evaluating the ionization state and gas cooling at each temperature independently.
@@ 2.1 Photoionization equilibrium (PIE)
When comparing model predictions with absorption line measurements, the gas is typically assumed to be in thermal equilibrium. Figure 2 depicts the cooling and heating rates for gas at solar and sub-solar metallicities, assuming a fiducial gas density of . Note that the metallicity for the model curves in Figure 2 refers to the total gas-phase metallicity, assuming a solar abundance pattern between elements. The equilibrium temperature depends on the gas metallicity, with higher enrichment levels leading to lower temperatures. The equilibrium temperature corresponding to the best-fit ionization model can be compared with the temperature estimated by comparing the absorption line widths from ionic species with sufficiently different masses.
2.2 Time-dependent photoionization (TDP)
@@
In equilibrium conditions, ion fractions () are assumed to be in a steady state at all times () because the gas cools slowly. However, chemically enriched gas clouds are expected to experience rapid cooling through collisional, free-bound, and free-free interactions. Therefore, the gas becomes "overionized" relative to the equilibrium scenario and a steady state no longer holds. In such cases, equations of ionization balance must be solved simultaneously with energy conservation to specify over time.
@@
Non-equilibrium collisional ionization models from Gnat & Sternberg (2007) have been compared with absorption line systems arising from hot gas (see e.g., Tripp et al., 2011; Meiring et al., 2013; Hussain et al., 2015; Nevalainen et al., 2017; Rosenwasser et al., 2018). However, photoionization from the extragalactic UVB cannot be neglected at temperatures relevant for some high-ionization species. The inclusion of a radiation background in time-dependent calculations modifies the ionization state of gas through photoionization and photoheating (see e.g., Oppenheimer & Schaye, 2013; Gnat, 2017).
@@
TDP calculations for this work are performed in CLOUDY following the hazy1 documentation, which provides a non-equilibrium cooling calculation script for modeling the radiative time-dependent cooling in a plane parallel gas slab without any ionizing radiation (see e.g., Gnat & Sternberg, 2007). The default script is modified to include ionizing radiation from a UVB 111See this ReadMe for a tutorial on building a suite of TDP models.. Cooling is assumed to be isochoric, and the calculations stop at thermal equilibrium or , whichever condition is achieved first. The gas is also assumed to be optically thin for these calculations.
2.3 Comparisons between PIE and TDP
A comparison between PIE (§ 2.1) and TDP ion fractions is illustrated in the left panel of Figure 3 for three commonly adopted tracers of warm-hot ionized gas, C IV, N V, and O VI, in QSO absorption-line studies. Note that the PIE fractions vary with gas density but do not depend on metallicity, while the predicted TDP fractions have a simultaneous density and metallicity dependence. The discrepancy between TDP and PIE is therefore metallicity dependent. At high temperatures, the metallicity-dependence is weaker under TDP because free-free cooling dominates. For a solar metallicity gas of density , O VI can be overabundant by an order of magnitude at . N V behaves similarly as O VI, but C IV is underabundant relative to PIE at this density because of an overabundance in C V at these cooler temperatures (see also Oppenheimer & Schaye, 2013; Gnat, 2017, for a comparison).
@@ 2.4 Model assumptions
While TDP models provide a self-consistent framework for examining metal-enriched absorption systems, their underlying assumptions can have a notable effect on the expected ion fractions. Three key ingredients of time-dependent photoionization models are the assumed ionizing radiation, the initial temperature from which cooling occurs, and the adopted abundance pattern. Reasonable choices for each ingredient are explored in the subsections here.
@@ 2.4.1 UVB prescription
The right panel in Figure 4 shows the predicted TDP fractions for C IV, N V, and O VI for three different UVBs at , including HM12, KS19, and FG20. The predictions are obtained using the procedure described in § 2.2. The high-temperature () peak comes from collisional ionization, while the low-temperature enhancement occurs because of recombination lags arising from rapid gas cooling. The fiducial choice of and for the density and metallicity are relevant to enriched C IV absorbers. While ion fractions for C IV are robust against UVB choices, N V and O VI TDP fractions can differ by a factor of up to three at . To facilitate a more focused discussion on the effect of TDP, FG20 is adopted as the fiducial UVB spectrum in the subsequent discussions.